Does any solid silver chromate form when 2.7x10-5 g of silver nitrate is dissolved in 15.0ml of 4.0x10-4 M potassium chromate?
Answer #1

Number of moles of AgNO3 dissolved= 2.7 ×10-5g/ 169.88g/mol = 1.59×10-7mol

number of Ag+ dissolved = 1.59×10-7mol

Volume of solution = 15ml

[Ag+] = (1.59 ×10-7mol/15ml) ×1000ml = 1.06×10-5M

[CrO42-] = 4.0 ×10-4M

Ag2CrO4(s) <--------> 2Ag+(aq) + CrO42-(aq)

Ksp = [Ag+]2[CrO42-] = 1.10 ×10-12

Qsp = [Ag+]2[CrO42-]

Qsp = (1.06 ×10-5M )2 × 4.0 ×10-4M

Qsp = 4.49 × 10-14

Qsp < Ksp

Therefore

Solid silver chromate would not form